# Reduced fidelities for free fermions out of equilibrium: From dynamical quantum phase transitions to Mpemba effect

### v1

@@ Abstract
We investigate the out-of-equilibrium dynamics after a quantum quench of the reduced fidelities between the states of a subregion $A$ at different times. Precisely, we consider the fidelity between the time-dependent state of $A$ and its initial value, as well as with the state at infinite time. We denote these fidelities as the reduced Loschmidt echo (RLE) and the final-state fidelity (FSF), respectively. If region $A$ is the full systems, the RLE coincide with the standard Loschmidt echo. We focus on quenches from Gaussian states in several instances of the XY spin chain. In the hydrodynamic limit of long times and large sizes of $A$, with their ratio fixed, the reduced fidelities admit a quasiparticle picture interpretation. Interestingly, for some quenches in the hydrodynamic regime the RLE features a complicated structure with an infinite sequence of nested lightcones, corresponding to quasiparticles with arbitrary large group velocities. This leads to a “staircase” of cusp-like singularities in the time-derivative of the fidelity. At the sub-hydrodynamic regime for some quenches the RLE exhibits cusp-like singularities, similar to the so-called dynamical quantum phase transitions (DQPT). We conjecture a criterion for the occurrence of the DQPT and for the “critical” times at which the singularities occur. Finally, we discuss the hydrodynamic limit of the FSF. In particular, we show that it provides a valuable tool to detect the so-called quantum Mpemba effect.
@@ 1 Introduction
Going beyond the revival probability, the analytic structure of the rate function (2) also contains valuable insights in the many-body dynamics. Indeed, for certain quantum systems in the thermodynamic limit $N\to\infty$, the rate function exhibits non-analyticities as a function of time. These phenomena are known as dynamical quantum phase transitions (DQPT) [1, 48, 49]. Generically, DQPTs are expected to occur when the initial state and the ground state of the quenched Hamiltonian belong to different quantum phases However, there are numerous exceptions to this behavior, and DQPT should be seen as a singular behavior distinct from equilibrium criticality [49].
@@ 5.1 Quasiparticle picture
where $\hat{\tilde{J}}_{x,k}(t)$ is constructed from the product $\hat{J}_{k}(t)\hat{J}_{k}(\infty)$, with $\hat{J}_{k}(\infty)$ obtained from $\hat{J}_{k}(t)$ by removing the time-dependent off-diagonal terms, which vanish via dephasing in the long-time limit. Moreover, in $\hat{\tilde{J}}_{x,k}$ one has to set the time-dependent phases to $1$, which equivalent to consider $\hat{\tilde{J}}_{x,k}(t=0)$. The time dependence in $\hat{\tilde{J}}_{x,k}(t)$ is taken into account by the quasiparticle propagation. Similarly as in Eq. (92), the superscript $\mathcal{A}$ in Eq. (93) denotes the restriction to rows and columns that correspond to the species in $A$. Now, we should observe that since $\hat{J}_{k}(\infty)$ is diagonal, the contributions that correspond to only one species in $A$ cancel out with the denominator in the definition of the FSF, see Eq. (20). This implies that the FSF is nonzero only if both the entangled pairs are in $A$. By employing Eq. (92) for the denominators and Eq. (93) for the numerator in the definition of the FSF, one recovers Eqs. (89), (90) and (91), after using that the FSF is real.
@@
5.2 Quantum Mpemba effect
@@
Let us consider a nonequilibrium situation where the Hamiltonian $H$ generating the time evolution possesses a certain symmetry. For simplicity, we assume a $U(1)$ symmetry with charge $Q$, namely $[H,Q]=0$. We thus expect that, irrespective of the initial state, all stationary states also possess that symmetry, and thus satisfy $\rho_{A}^{\infty},Q_{A}]=0$, where $Q_{A}$ is the charge restricted to subsystem $A$. Given an initial state $|\Psi_{0}\rangle$ which breaks the symmetry, $|\Psi_{0}\rangle\langle\Psi_{0}|,Q]\neq 0$, a natural question is to quantify how fast is the symmetry restored during the time evolution, or how fast does the system thermalize and reaches its stationary state. Certain situations give rise to the quantum Mpemba effect, where dynamics from initial states which strongly break the symmetry restore it faster than dynamics from initial states closer to a symmetric state [4, 5]. The quantum Mpemba effect has been investigated theoretically and experimentally using various measures, such as the so-called entanglement asymmetry[4, 67, 68], the trace distance [69, 70], the Frobenuis distance [71] or the relative entropy [72, 73] (see Ref. [5] for a recent review on the subject).
@@
By design, the FSF is a powerful tool to probe thermalization in quantum systems out of equilibrium, and in particular the rate at which the system reaches its stationary state after a quench. Moreover, reduced fidelities have a strong experimental relevance, as we already discussed for the RLE. We thus expect the logarithmic FSF to be a natural quantity for investigating quantum Mpemba effects, both analytically and experimentally. However, as opposed to the entanglement asymmetry which directly quantifies symmetry breaking and its restoration, the RLE is an indicator of the relaxation towards a symmetric stationary state. These two quantities are thus complementary in the study of thermalization in quantum systems. To test this claim, we consider quenches in the XX chain ($h=\gamma=0$) from ground states of the XY chain with various values of $h_{0},\gamma_{0}$. The XX Hamiltonian possesses a $U(1)$ symmetry, where the associated charge is simply the fermion number, whereas ground states of the XY chain with arbitrary anisotropy $\gamma_{0}$ break this symmetry. Quantum Mpemba effects, or lack thereof, have been diagnosed in this quench protocol via the entanglement asymmetry [68]. In Figs. 9 and 10, we consider the same quenches as in Ref. [68], and obtain exactly the same qualitative results. In particular, in Fig. 9 the crossings indicate that certain states initially further away from their stationary values thermalize faster than other more symmetric initial states, hence corresponding to a quantum Mpemba effect. In contrast, in Fig. 10 we display situations where such crossings do not appear, and where more symmetric initial states do thermalize faster than less symmetric ones.
@@
To understand these results quantitatively, it is important to investigate under which conditions Mpemba effects occurs. Let us consider two quenches $\mathcal{Q}_{1}$ and $\mathcal{Q}_{2}$ in the XX chain, where $\mathcal{Q}_{j}=(h_{0}^{(j)},\gamma_{0}^{(j)})$ encodes the data of the initial states, and $\mathcal{Q}_{1}\neq\mathcal{Q}_{2}$. We assume without loss of generality that initially we have $\Lambda_{A}^{\infty,\mathcal{Q}_{1}}(0)>\Lambda_{A}^{\infty,\mathcal{Q}_{2}}(0)$. By definition, a quantum Mpemba effect occurs if there is a crossing time $t_{C}$ such that for all $t>t_{C}$ the inequality is reversed, $\Lambda_{A}^{\infty,\mathcal{Q}_{1}}(t)<\Lambda_{A}^{\infty,\mathcal{Q}_{2}}(t)$. Since the FSF is a monotonic decreasing quantity, the condition for a quantum Mpemba effect reduces to the existence of a finite time $0<t_{C}<\infty$ such that
@@
and this condition is both necessary and sufficient. There is no explicit formula for the crossing time $t_{C}$, but the exact QPP form of $\Lambda_{A}^{\infty}(t)$ (see Eq. (97)) allows us to derive it precisely with Eq. (98).
@@
where $\hat{k}=v^{-1}(\ell/(2t))$, and $v^{-1}(x)=\arcsin(x)$ in the XX quench under consideration here. The first condition in Eq. (99) does not depend on time and simply reflects the initial condition $\Lambda_{A}^{\infty,\mathcal{Q}_{1}}(0)>\Lambda_{A}^{\infty,\mathcal{Q}_{2}}(0)$. The second, however, involves the modes with momenta in $\hat{k},\hat{k}$], which are the pairs of quasiparticles that are in $A$ at time $t$. Moreover, $\Lambda_{k}^{\mathcal{Q}_{j}}$ and $\Lambda_{k+\pi}^{\mathcal{Q}_{j}}$ are the contributions to the FSF due to the members of the generic entangled pair. Finally, since we are in the large-time limit, we only consider the slowest modes and hence neglect the contributions proportional to the velocity term $2|v_{k}|t$ in the integrals. Let us also stress that here, the large-time assumption is $\ell/(2t)<1$, otherwise the momenta $\hat{k}$ is not well defined. In terms of the QPP interpretation, this means that we wait at least for the fastest moving quasiparticles initially created in the center of $A$ to leave the subsystem, after a time $t=v_{\rm max}\ell/2=\ell/2$. Typically, we thus have $t_{I}>t_{C}$, because the Mpemba crossing occurring at time $t_{C}$ can occur for $t_{C}\ell<1/2$, and hence before the validity of the large-time criterion given in Eq. (99). For instance, (i) in the left panel of Fig. 9 for $\mathcal{Q}_{1}=(0.5,0.6)$ (green curve) and $\mathcal{Q}_{2}=(1.1,0.6)$ (orange curve) we have $t_{C}/\ell=0.48$, (ii) in the right panel of Fig. 9 for $\mathcal{Q}_{1}=(-0.1,0.7)$ (orange curve) and $\mathcal{Q}_{2}=(0.8,0.4)$ (pink curve) we have $t_{C}/\ell=0.37$, and (iii) in the same panel for $\mathcal{Q}_{1}=(-0.1,1.2)$ (green curve) and $\mathcal{Q}_{2}=(0.8,0.4)$ (pink curve) we have $t_{C}/\ell=0.48$.

### v2

@@ Abstract
We investigate the out-of-equilibrium dynamics after a quantum quench of the reduced fidelities between the states of a subregion $A$ at different times. Precisely, we consider the fidelity between the time-dependent state of $A$ and its initial value, as well as with the state at infinite time. We denote these fidelities as the reduced Loschmidt echo (RLE) and the final-state fidelity (FSF), respectively. If region $A$ is the full system, the RLE coincides with the standard Loschmidt echo. We focus on quenches from Gaussian states in several instances of the XY spin chain. In the hydrodynamic limit of long times and large sizes of $A$, with their ratio fixed, the reduced fidelities admit a quasiparticle picture interpretation. Interestingly, for some quenches in the hydrodynamic regime the RLE features a complicated structure with an infinite sequence of nested lightcones, corresponding to quasiparticles with arbitrary large group velocities. This leads to a “staircase” of cusp-like singularities in the time-derivative of the fidelity. At the sub-hydrodynamic regime for some quenches the RLE exhibits cusp-like singularities, similar to the so-called dynamical quantum phase transitions (DQPT). We conjecture a criterion for the occurrence of the DQPT and for the “critical” times at which the singularities occur. Finally, we discuss the hydrodynamic limit of the FSF. In particular, we show that it provides a valuable tool to detect the so-called quantum Mpemba effect.
@@ 1 Introduction
Going beyond the revival probability, the analytic structure of the rate function (2) also contains valuable insights in the many-body dynamics. Indeed, for certain quantum systems in the thermodynamic limit $N\to\infty$, the rate function exhibits non-analyticities as a function of time. These phenomena are known as dynamical quantum phase transitions (DQPT) [1, 48, 49]. Generically, DQPTs are expected to occur when the initial state and the ground state of the quenched Hamiltonian belong to different quantum phases. However, there are numerous exceptions to this behavior, and DQPT should be seen as a singular behavior distinct from equilibrium criticality [49].
@@ 5.1 Quasiparticle picture
where $\hat{\tilde{J}}_{x,k}(t)$ is constructed from the product $\hat{J}_{k}(t)\hat{J}_{k}(\infty)$, with $\hat{J}_{k}(\infty)$ obtained from $\hat{J}_{k}(t)$ by removing the time-dependent off-diagonal terms, which vanish via dephasing in the long-time limit. Moreover, in $\hat{\tilde{J}}_{x,k}$ one has to set the time-dependent phases to $1$, which is equivalent to consider $\hat{\tilde{J}}_{x,k}(t=0)$. The time dependence in $\hat{\tilde{J}}_{x,k}(t)$ is taken into account by the quasiparticle propagation. Similarly as in Eq. (92), the superscript $\mathcal{A}$ in Eq. (93) denotes the restriction to rows and columns that correspond to the species in $A$. Now, we should observe that since $\hat{J}_{k}(\infty)$ is diagonal, the contributions that correspond to only one species in $A$ cancel out with the denominator in the definition of the FSF, see Eq. (20). This implies that the FSF is nonzero only if both the entangled pairs are in $A$. By employing Eq. (92) for the denominators and Eq. (93) for the numerator in the definition of the FSF, one recovers Eqs. (89), (90) and (91), after using that the FSF is real.
@@
5.2 Quantum Mpemba effect with symmetry restoration
@@
Let us consider a nonequilibrium situation where the Hamiltonian $H$ generating the time evolution possesses a certain symmetry. For simplicity, we assume a $U(1)$ symmetry with charge $Q$, namely $[H,Q]=0$. We thus expect that, irrespective of the initial state, all stationary states also possess that symmetry, and thus satisfy $\rho_{A}^{\infty},Q_{A}]=0$, where $Q_{A}$ is the charge restricted to subsystem $A$. Given an initial state $|\Psi_{0}\rangle$ which breaks the symmetry, $|\Psi_{0}\rangle\langle\Psi_{0}|,Q]\neq 0$, a natural question is to quantify how fast is the symmetry restored during the time evolution, or how fast does the system thermalize and reaches its stationary state. Certain situations give rise to the quantum Mpemba effect, where dynamics from initial states which strongly break the symmetry restore it faster than dynamics from initial states closer to a symmetric state [4, 5]. The quantum Mpemba effect has been investigated theoretically and experimentally using various measures, such as the entanglement asymmetry[4, 67, 68], the trace distance [69, 70], the Frobenius distance [71] or the relative entropy [72, 73]. Quantum fidelities akin to the FSF have also been used to detect quantum Mpemba effects in two dimensional bosonic systems [74]. For a recent review on quantum Mpemba effects, see Ref. [5].
@@
By design, the FSF is a powerful tool to probe thermalization in quantum systems out of equilibrium, and in particular the rate at which the system reaches its stationary state after a quench. Moreover, reduced fidelities have a strong experimental relevance, as we already discussed for the RLE. We thus expect the logarithmic FSF to be a natural quantity for investigating quantum Mpemba effects, both analytically and experimentally. However, as opposed to the entanglement asymmetry which directly quantifies symmetry breaking and its restoration, the FSF is an indicator of the relaxation towards a symmetric stationary state. These two quantities are thus complementary in the study of thermalization in quantum systems. To test this claim, we consider quenches in the XX chain ($h=\gamma=0$) from ground states of the XY chain with various values of $h_{0},\gamma_{0}$. The XX Hamiltonian possesses a $U(1)$ symmetry, where the associated charge is simply the fermion number, whereas ground states of the XY chain with arbitrary anisotropy $\gamma_{0}$ break this symmetry. Quantum Mpemba effects, or lack thereof, have been diagnosed in this quench protocol via the entanglement asymmetry [68]. In Figs. 9 and 10, we consider the same quenches as in Ref. [68], and obtain exactly the same qualitative results. In particular, in Fig. 9 the crossings indicate that certain states initially further away from their stationary values thermalize faster than other more symmetric initial states, hence corresponding to a quantum Mpemba effect. In contrast, in Fig. 10 we display situations where such crossings do not appear, and where more symmetric initial states do thermalize faster than less symmetric ones.
@@
5.3 Quantum Mpemba effect without symmetry restoration
Mpemba effects are not always related to a symmetry restoration. Indeed, there are quench protocols where dynamics from initial states further away from the stationary state relax faster, even though the Hamiltonian driving the time evolution does not have symmetries which are broken by the initial states. Such Mpemba effects were for instance observed in the Ising quench using the relative entropy between the time-evolved RDM and the stationary state [73].
As already mentioned, the FSF is a probe of how fast a system relaxes after a quench. In particular, it is still well-defined and relevant even in contexts where there is no symmetry restoration in the dynamics, as opposed to the entanglement asymmetry. We thus expect it to be able to diagnose Mpemba effects, even without symmetry restoration. To test this, we consider the same quench protocols as in Ref. [73]: a quench to the Ising chain with $\gamma=1$ and $h=0.2$ from the Ising chain ($\gamma_{0}=1$) with various values of $h_{0}$. We illustrate our results in Fig. 11. As expected, we observe the same qualitative crossings (or lack thereof) as in Ref. [73]. Moreover, we again find a perfect agreement between the numerical results and the QPP formula (97) for the FSF.
Figure 11: Logarithmic FSF in the Ising chain ($\gamma=1$) with $h=0.2$ and $\ell=200$. The initial states are ground states of the Ising chain ($\gamma_{0}=1$) with various values of $h_{0}$. The symbols are obtained by exact numerical diagonalization, and the solid lines are the analytical prediction of Eq. (91).
5.4 Conditions for Mpemba effects
To understand our results quantitatively, it is important to investigate under which conditions Mpemba effects occurs. Let us consider two quenches $\mathcal{Q}_{1}$ and $\mathcal{Q}_{2}$ where $\mathcal{Q}_{j}=(h_{0}^{(j)},\gamma_{0}^{(j)},h^{(j)},\gamma^{(j)})$ encodes the data of the quench $j$, and $\mathcal{Q}_{1}\neq\mathcal{Q}_{2}$. We assume without loss of generality that initially we have $\Lambda_{A}^{\infty,\mathcal{Q}_{1}}(0)>\Lambda_{A}^{\infty,\mathcal{Q}_{2}}(0)$. By definition, a quantum Mpemba effect occurs if there is a crossing time $t_{C}$ such that for all $t>t_{C}$ the inequality is reversed, $\Lambda_{A}^{\infty,\mathcal{Q}_{1}}(t)<\Lambda_{A}^{\infty,\mathcal{Q}_{2}}(t)$. The QPP formula (97) for the FSF is monotonically decreasing. Therefore, in the thermodynamic limit, the condition for a quantum Mpemba effect reduces to the existence of a finite time $0<t_{C}<\infty$ such that
@@ 5.4 Conditions for Mpemba effects
and this condition is typically both necessary and sufficient. There is no explicit formula for the crossing time $t_{C}$, but the exact QPP form of $\Lambda_{A}^{\infty}(t)$ (see Eq. (97)) allows us to derive it precisely with Eq. (98). We stress that there are examples of quenches with multiple crossing [75], where different times $t_{C}^{(i)}$ satisfy Eq. (98). In this case there is an Mpemba effect if the number of crossing times is odd.
@@
where $\hat{k}=v^{-1}(\ell/(2t))$. For instance, $v^{-1}(x)=\arcsin(x)$ in the XX quench. The first condition in Eq. (99) does not depend on time and simply reflects the initial condition $\Lambda_{A}^{\infty,\mathcal{Q}_{1}}(0)>\Lambda_{A}^{\infty,\mathcal{Q}_{2}}(0)$. The second, however, involves the modes with momenta in $\hat{k},\hat{k}$], which are the pairs of quasiparticles that are in $A$ at time $t$. Moreover, $\Lambda_{k}^{\mathcal{Q}_{j}}$ and $\Lambda_{k+\pi}^{\mathcal{Q}_{j}}$ are the contributions to the FSF due to the members of the generic entangled pair. Finally, since we are in the large-time limit, we only consider the slowest modes and hence neglect the contributions proportional to the velocity term $2|v_{k}|t$ in the integrals. Let us also stress that for the XX quench, the large-time assumption is $\ell/(2t)<1$, otherwise the momenta $\hat{k}$ is not well defined. In terms of the QPP interpretation, this means that we wait at least for the fastest moving quasiparticles initially created in the center of $A$ to leave the subsystem, after a time $t=v_{\rm max}\ell/2$, with $v_{\rm max}=1$ for the XX quench. Typically, we thus have $t_{I}>t_{C}$, because the Mpemba crossing occurring at time $t_{C}$ can occur for $t_{C}/\ell<v_{\rm max}/2$, and hence before the validity of the large-time criterion given in Eq. (99). For instance, (i) in the left panel of Fig. 9 for $\mathcal{Q}_{1}=(0.5,0.6)$ (green curve) and $\mathcal{Q}_{2}=(1.1,0.6)$ (orange curve) we have $t_{C}/\ell=0.48$, (ii) in the right panel of Fig. 9 for $\mathcal{Q}_{1}=(-0.1,0.7)$ (orange curve) and $\mathcal{Q}_{2}=(0.8,0.4)$ (pink curve) we have $t_{C}/\ell=0.37$, and (iii) in the same panel for $\mathcal{Q}_{1}=(-0.1,1.2)$ (green curve) and $\mathcal{Q}_{2}=(0.8,0.4)$ (pink curve) we have $t_{C}/\ell=0.48$.